# Two-component atomic Fermi superfluid with spin-orbital coupling in thin spherical-shell geometry

### v1

@@ I Introduction
We will present several findings of the influence of SOC on the spherical Fermi superfluid in BCS-BEC crossover: (1) The chemical potential on the BCS side is suppressed by SOC, a phenomenon also observable in Fermi gases without pairing interaction subject to the same SOC. (2) The order parameter of the Fermi superfluid no longer determines the energy gap of the energy dispersion in the presence of SOC. We will present numerical evidence where the minimum of the energy dispersion approaches zero while the order parameter remains finite. (3) By evaluating the isothermal compressibility, we observe kinks as the pairing interaction or SOC coupling constant changes. Those kinks hint singular behavior of higher derivatives of the free energy. (4) At the compressibility kinks, the order parameter remains finite and indicates the superfluid survives while the minimum of the energy dispersion plunges toward zero. In contrast, the compressibility of a Fermi gas with SOC but without pairing interaction does not exhibit any kink. (5) The aforementioned phenomena survive when the SOC deviates from the isotropic form and becomes anisotropic.
The rest of the paper is organized as follows. Sec. II presents the theory of two-component atomic Fermi superfluid with SOC across the BCS-BEC crossover on a spherical surface. In addition to the equations of state determining the order parameter and chemical potential, the isothermal compressibility is derived for later analyses. Sec. III shows the numerical results of the spherical Fermi superfluid with SOC. The decoupling of the energy gap from the order parameter will be clearly demonstrated. The compressibility exhibits kinks indicating singular behavior of higher derivatives of the free energy, which occur whenever the energy gap approaches zero. Sec. IV shows that similar phenomena are observable when the SOC becomes anisotropic. Sec. V discusses possible realizations and measurements of the predictions and their implications. Finally, Sec. VI concludes our work. The Appendix summarizes the corresponding behavior of a two-component Fermi gas without pairing interaction but subject to the same SOC and spherical-shell geometry to contrast the differences between spherical normal gas and superfluid in the presence of SOC.
@@ II.3 Isothermal compressibility
$\frac{\partial^{2}E_{l}}{\partial\mu^{2}}=\frac{\Delta^{2}}{E_{l}^{3}},\quad\frac{\partial^{2}E_{l}}{\partial\mu\partial\Delta}=\frac{\xi_{l}\Delta}{E_{l}^{3}},\quad\frac{\partial^{2}E_{l}}{\partial\Delta^{2}}=\frac{1}{E_{l}}-\frac{\Delta^{2}}{E_{l}^{3}}.$ (32)
@@ III.1 BCS-BEC crossover with SOC on sphere
Figure 1: (a) Chemical potential $\mu$ and (b) superfluid order parameter $\Delta$ as functions of $-\ln(k_{F}a)$ on a spherical shell with SOC at $T=0$. All quantities have been normalized. The black, red, and blue lines correspond to $\gamma=0$, $10$ and $20$, respectively. Here $E_{F}/E_{0}=110$.
@@ III.2 Isothermal compressibility and singular behavior
In the absence of SOC, $E_{min}=\Delta$ if $\mu>0$ and $E_{min}=\sqrt{\mu^{2}+\Delta^{2}}$ if $\mu<0$ for Fermi superfluid in the BCS-BEC crossover [55]. However, one can see that the energy gap given by $E_{min}$ almost vanishes in the presence of SOC on the BCS side but grows on the BEC side back to the value without SOC. However, $E_{min}$ is never truly zero because of avoided crossing due to the finite order parameter $\Delta$. This is in contrast to the Fermi gas with the same SOC but no pairing interaction shown in the Appendix, where the energy dispersion crosses zero with no energy gap in the absence of $\Delta$.
@@
In Fig. 3, we plot $\mu$ and $\Delta$ as functions of $\gamma$ for a case on the BCS side with $-\ln(k_{F}a)=-1.5$ (black lines) and one in the middle of the crossover with $-\ln(k_{F}a)=-0.5$ (red lines). In panel (a), one can see that $\mu$ decreases from positive to negative in both cases as $\gamma$ increases. Moreover, $\mu$ starts with positive values when $\gamma=0$ but changes signs around $\gamma\approx\sqrt{N}$, which is consistent with the Fermi gas without pairing interaction discussed in the Appendix. In panel (b), one can see that $\Delta$ is suppressed by increasing $\gamma$ in both cases, which is consistent with the suppression of $\Delta$ shown in Fig. 1(b). In the deep-BCS case with $-\ln(k_{F}a)=-1.5$, $\Delta$ approaches zero as $\gamma$ increases but still remains finite. Therefore, the Fermi gas is still a superfluid in its ground state.
@@
Fig. 4 shows the energy gap $E_{min}$ from the energy dispersions and $\kappa_{T}$ for the same cases shown in Fig. 3. We remark that in the presence of SOC, the energy gap is no longer determined by the order parameter $\Delta$, and we numerically scan $E_{1,2}$ to find the minimal value as $E_{min}$. One can see in panel (a) that there are a few places where the energy gap becomes extremely small. However, the presence of avoided crossing due to the small but non-zero $\Delta$ manages to keep the energy gap above zero. As shown in panel (b), there are three peaks of $\kappa_{T}$ in the deep BCS case shown by the solid curve. The singular behavior of $\kappa_{T}$ indicates singular higher-derivatives of the free energy as $\gamma$ changes. However, $\Delta$ does not vanish, as shown in Fig. 3, thereby showcasing the non-conventional behavior of a Fermi superfluid on a sphere induced by SOC. We caution that the compressibility peaks in panel (b) result from a complex interplay between the change of the energy dispersions with SOC, superfluidity, and the spherical geometry. To verify this, the curve of $\kappa_{T}$ of a Fermi gas with the same SOC but no pairing interaction is shown by the blue dotted line in panel (b), which exhibits smooth behavior without any kinks.
@@
To investigate the origin of the singular higher derivatives of the free energy indicated by the kinks of $\kappa_{T}$ shown in Fig. 4(b), we analyze the energy dispersions $E_{1,2}$ on the two sides of the highest peak in Fig. 4(b) located at $\gamma\approx 15$. Explicitly, Fig. 5 shows the energy dispersions for $\gamma=14$ and $15.5$, respectively. One can see that the dispersions indeed exhibit different structures: The lower branch approaches zero twice away from $l=0$ when $\gamma=14$ but only once away from $l=0$ when $\gamma=15.5$. The insets of Fig. 5 emphasize that the energy gap can be extremely small yet still non-zero. Therefore, the structural change in energy space contributes to the suppressed minimum of the energy spectrum, which occurs when $\kappa_{T}$ exhibits a kink. However, the presence of a small but finite order parameter $\Delta$ plays a significant role in adjusting the energy dispersions and thermodynamic quantities. In Fig. 7 in the Appendix, we show that a two-component Fermi gas without pairing interaction but with the same SOC may exhibit structural changes in its energy dispersions but nevertheless has smooth $\kappa_{T}$, indicating no singularity in its free energy.
@@
We remark that in our numerical calculations, a cutoff of $L_{\text{cutoff}}=30L_{m}$ is introduced when carrying out the summations, where $L_{m}$ is the highest occupied level for a corresponding noninteracting two-component Fermi gas with the same total particle number. As long as the cutoff energy is much larger than the Fermi energy, the numerical results are insensitive to slight variations of the cutoff. This is indeed the case observed in our results, and we have checked that $\mu$ and $\Delta$ are basically the same if $L_{\text{cutoff}}$ is further increased. Moreover, the energy minimum occurs when $l$ is relatively small, as shown in Fig. 5. Therefore, increasing the cutoff $L_{\text{cutoff}}$ does not change the location and magnitude of $E_{min}$ and leaves the peaks of $\kappa_{T}$ intact as well. We also mention that due to the particle-hole mixing in the superfluid when $\Delta$ is finite, although the dispersions of $E_{1,2}$ may look similar to the absolute values of the dispersion $|E_{1,2}^{0}|$ of the Fermi gas without pairing interaction discussed in the Appendix, the minimum of the dispersion of the Fermi superfluid does not reach zero due to avoided crossing even if $\Delta$ is suppressed by the presence of SOC.
@@ IV Anisotropic SOC
Figure 6: (a) Order parameter $\Delta$ and (b) isothermal compressibility $\kappa_{T}$ of Fermi superfluid with the anisotropic SOC terms shown in Eq. (41) as functions of $\gamma_{3}$ on a spherical surface at $T=0$. Here $\gamma_{1}=\gamma_{2}=\gamma_{3}/2$ and $-\ln(k_{F}a)=-1.5$. Here $\kappa_{F}$ is the compressibility of a Fermi gas with the same SOC and total particle number but no pairing interaction.
@@ V Implications
Since atomic BEC in a thin spherical shell has been realized using bubble traps or multi-species phase separation [1, 2, 3, 4, 5], similar techniques may apply to fermionic atoms to investigate Fermi gases on a spherical surface. As explained in Ref. [25], it may take more internal states and confining potentials to trap two-component Fermi gases in a thin spherical shell. After introducing tunable attractive interactions between the fermionic atoms via Feshbach resonance [56, 57], the ground state is expected to be a Fermi superfluid due to the formation of Cooper pairs.
@@
We mention that the presence of SOC can lead to very different behavior between Fermi gases with and without pairing interaction in a thin spherical shell by contrasting the aforementioned results and those in the Appendix. Instead of destroying the superfluid, SOC modifies the particle-hole mixed excitation spectrum and causes singular behavior in the free energy, which can be inferred by the kinks of the compressibility. Even without SOC, the compressibility of Fermi superfluid in 3D drastically differs from that on a spherical surface due to different dimensionality and geometry. Therefore, atomic Fermi superfluid in a thin spherical shell provides a playground for integrating superfluid, SOC, and geometric effects.
@@ Acknowledgements.
Y. H. was supported by the National Key R&\&&D Program of China (No. 2024YFF0508503). C. C. C. was partly supported by the NSF (No. PHY-2310656).
@@ Appendix A Fermi gases with only SOC on sphere
Therefore, If $\gamma^{2}>N$, there is no positive $\mu$ solution.
For $\mu<0$, on the other hand, we always have $E_{1}^{(0)}>0$. The positive roots of equation $E_{2}^{(0)}=0$ are $L^{(2)}_{1,2}=\dfrac{\gamma\pm\sqrt{\gamma^{2}+4\mu}}{2}$. This situation is shown in the Bottom of Fig. 7. Thus we find that
@@
Consequently, we find a solution $\mu=\dfrac{N^{2}-\gamma^{4}}{4\gamma^{2}}<0$ when $\gamma^{2}>N$. In summary, when $\gamma^{2}>N$, we have $\mu<0$ for the Fermi gas with only SOC. In the BCS limit, there will be a small $\Delta$, but we will still expect that $\mu<0$ in this case.
@@
The corresponding compressibility is shown by the dotted line in Fig. 4. There is no singular behavior without pairing interaction as the strength of SOC varies. The difference between the Fermi gases with and without pairing interaction shows the importance of particle-hole mixing in the energy dispersions, which gives rise to interesting phenomena in superfluid only.

### v2

@@ I Introduction
We will present several findings of the influence of SOC on the spherical Fermi superfluid in BCS-BEC crossover: (1) The chemical potential on the BCS side is suppressed by SOC, a phenomenon also observable in Fermi gases without pairing interaction subject to the same SOC. (2) The order parameter of the Fermi superfluid no longer determines the energy gap of the energy dispersion in the presence of SOC. We will present numerical evidence where the minimum of the energy dispersion approaches zero while the order parameter remains finite. (3) By evaluating the isothermal compressibility, we observe kinks as the pairing interaction or SOC coupling constant changes. Those kinks hint singular behavior of higher derivatives of the free energy. (4) At the compressibility kinks, the order parameter remains finite and indicates the superfluid survives while the minimum of the energy dispersion plunges toward zero. In contrast, the compressibility of a Fermi gas with SOC but without pairing interaction or a planar Fermi superfluid does not exhibit any kink. (5) The aforementioned phenomena survive when the SOC deviates from the isotropic form and becomes anisotropic.
The rest of the paper is organized as follows. Sec. II presents the theory of two-component atomic Fermi superfluid with SOC across the BCS-BEC crossover on a spherical surface. In addition to the equations of state determining the order parameter and chemical potential, the isothermal compressibility is derived for later analyses. Sec. III shows the numerical results of the spherical Fermi superfluid with SOC. The decoupling of the energy gap from the order parameter will be clearly demonstrated. The compressibility exhibits kinks indicating singular behavior of higher derivatives of the free energy, which occur whenever the energy gap approaches zero. Sec. IV shows that similar phenomena are observable when the SOC becomes anisotropic. Sec. V discusses possible realizations and measurements of the predictions and their implications. Finally, Sec. VI concludes our work. Appendix A summarizes the corresponding behavior of a two-component Fermi gas without pairing interaction but subject to the same SOC and spherical-shell geometry to contrast the differences between spherical normal gas and superfluid in the presence of SOC. Appendix B shows the results of Fermi superfluids with SOC on a two-dimensional (2D) plane to contrast the geometric effects.
@@ II.3 Isothermal compressibility
$\frac{\partial^{2}E_{l}}{\partial\mu^{2}}=\frac{\Delta^{2}}{E_{l}^{3}},\leavevmode\nobreak\ \frac{\partial^{2}E_{l}}{\partial\mu\partial\Delta}=\frac{\xi_{l}\Delta}{E_{l}^{3}},\leavevmode\nobreak\ \frac{\partial^{2}E_{l}}{\partial\Delta^{2}}=\frac{1}{E_{l}}-\frac{\Delta^{2}}{E_{l}^{3}}.$ (32)
@@ III.1 BCS-BEC crossover with SOC on sphere
Figure 1: (a) Chemical potential $\mu$ and (b) superfluid order parameter $\Delta$ as functions of $-\ln(k_{F}a)$ on a spherical shell with SOC at $T=0$. All quantities have been normalized. The black, red, and blue lines correspond to $\gamma=0$, $10$, and $20$, respectively. Here $E_{F}/E_{0}=110$.
@@ III.2 Isothermal compressibility and singular behavior
In the absence of SOC and in an Euclidean geometry, the parabolic energy dispersion $\xi_{k}=k^{2}/(2M)-\mu$ in $E=\sqrt{\xi_{k}^{2}+\Delta^{2}}$ results in $E_{min}=\Delta$ if $\mu>0$ and $E_{min}=\sqrt{\mu^{2}+\Delta^{2}}$ if $\mu<0$ for Fermi superfluids in the BCS-BEC crossover [55]. The presence of SOC and the curved space on a sphere alter the energy dispersion in the spectrum and shift the magnitude and location of $E_{min}$ in the Fermi superfluid, thereby causing the decoupling between the order parameter $\Delta$ and the energy gap $E_{min}$. Moreover, one can see that the energy gap given by $E_{min}$ almost vanishes in the presence of SOC on the BCS side but grows on the BEC side back to the value without SOC. However, $E_{min}$ is never truly zero because of avoided crossing due to the finite order parameter $\Delta$. This is in contrast to the Fermi gas with the same SOC but no pairing interaction shown in the Appendix LABEL:FS, where the energy dispersion crosses zero with no energy gap in the absence of $\Delta$. We also caution that the order parameter $\Delta$ characterizes the symmetry breaking of the internal U(1) symmetry of the Fermi superfluid due to formation of Cooper pairs. Meanwhile, the energy gap $E_{min}$ quantifies the excitation energy to form quasi-particles out of the Fermi superfluid. The decoupling of $E_{min}$ and $\Delta$ in the presence of SOC shows that while the U(1) symmetry remains broken in the Fermi superfluid, the change in the energy spectrum indicates a lower excitation threshold particularly in the BCS regime.
@@
In Fig. 3, we plot $\mu$ and $\Delta$ as functions of $\gamma$ for a case on the BCS side with $-\ln(k_{F}a)=-1.5$ (black lines) and one in the middle of the crossover with $-\ln(k_{F}a)=-0.5$ (red lines). In panel (a), one can see that $\mu$ decreases from positive to negative in both cases as $\gamma$ increases. Moreover, $\mu$ starts with positive values when $\gamma=0$ but changes signs around $\gamma\approx\sqrt{N}$, which is consistent with the Fermi gas without pairing interaction discussed in Appendix A. In panel (b), one can see that $\Delta$ is suppressed by increasing $\gamma$ in both cases, which is consistent with the suppression of $\Delta$ shown in Fig. 1(b). In the deep-BCS case with $-\ln(k_{F}a)=-1.5$, $\Delta$ approaches zero as $\gamma$ increases but still remains finite. Therefore, the Fermi gas is still a superfluid in its ground state.
@@
Fig. 4 shows the energy gap $E_{min}$ from the energy dispersions and $\kappa_{T}$ for the same cases shown in Fig. 3. We remark that in the presence of SOC, the energy gap is no longer determined by the order parameter $\Delta$, and we numerically scan $E_{1,2}$ to find the minimal value as $E_{min}$. One can see in panel (a) that there are a few places where the energy gap becomes extremely small. However, the presence of avoided crossing due to the small but non-zero $\Delta$ manages to keep the energy gap above zero. As shown in panel (b), there are three peaks of $\kappa_{T}$ in the deep BCS case shown by the solid curve. The singular behavior of $\kappa_{T}$ indicates singular higher-derivatives of the free energy as $\gamma$ changes. However, $\Delta$ does not vanish, as shown in Fig. 3, thereby showcasing the non-conventional behavior of a Fermi superfluid on a sphere induced by SOC. We caution that the compressibility peaks in panel (b) result from a complex change of the energy dispersions due to SOC, superfluidity, and spherical geometry. To verify this, the curve of $\kappa_{T}$ of a Fermi gas with the same SOC but no pairing interaction is shown by the blue dotted line in panel (b), which exhibits smooth behavior without any kinks. The singular behavior of $\kappa_{T}$ due to SOC and superfluidity on the compact spherical geometry may hint possible structural transitions in energy space without additional symmetry breaking other than the already broken U(1) symmetry due to Cooper-pair formation. Identifying possible energy-space structural transitions induced by SOC, superfluidity, and spherical geometry awaits future research.
@@
To investigate the origin of the singular higher derivatives of the free energy indicated by the kinks of $\kappa_{T}$ shown in Fig. 4(b), we analyze the energy dispersions $E_{1,2}$ on the two sides of the highest peak in Fig. 4(b) located at $\gamma\approx 15$. Explicitly, Fig. 5 shows the energy dispersions for $\gamma=14$ and $15.5$, respectively. One can see that the dispersions indeed exhibit different structures: The lower branch approaches zero twice away from $l=0$ when $\gamma=14$ but only once away from $l=0$ when $\gamma=15.5$. The insets of Fig. 5 emphasize that the energy gap can be extremely small yet still non-zero. Therefore, the structural change in energy space contributes to the suppressed minimum of the energy spectrum, which occurs when $\kappa_{T}$ exhibits a kink. However, the presence of a small but finite order parameter $\Delta$ plays a significant role in adjusting the energy dispersions and thermodynamic quantities. In Fig. 7 in Appendix A, we show that a two-component Fermi gas without pairing interaction but with the same SOC may exhibit structural changes in its energy dispersions but nevertheless has smooth $\kappa_{T}$, indicating no singularity in its free energy.
@@
We remark that in our numerical calculations, a cutoff of $L_{\text{cutoff}}=30L_{m}$ is introduced when carrying out the summations, where $L_{m}$ is the highest occupied level for a corresponding noninteracting two-component Fermi gas with the same total particle number. As long as the cutoff energy is much larger than the Fermi energy, the numerical results are insensitive to slight variations of the cutoff. This is indeed the case observed in our results, and we have checked that $\mu$ and $\Delta$ are basically the same if $L_{\text{cutoff}}$ is further increased. Moreover, the energy minimum occurs when $l$ is relatively small, as shown in Fig. 5. Therefore, increasing the cutoff $L_{\text{cutoff}}$ does not change the location and magnitude of $E_{min}$ and leaves the peaks of $\kappa_{T}$ intact as well. We also mention that due to the particle-hole mixing in the superfluid when $\Delta$ is finite, although the dispersions of $E_{1,2}$ may look similar to the absolute values of the dispersion $|E_{1,2}^{0}|$ of the Fermi gas without pairing interaction discussed in Appendix A, the minimum of the dispersion of the Fermi superfluid does not reach zero due to avoided crossing even if $\Delta$ is suppressed by the presence of SOC.
@@ IV Anisotropic SOC
Figure 6: (a) Order parameter $\Delta$ and (b) isothermal compressibility $\kappa_{T}$ of a Fermi superfluid with the anisotropic SOC term shown in Eq. (41) as functions of $\gamma_{3}$ on a spherical surface at $T=0$. Here $\gamma_{1}=\gamma_{2}=\gamma_{3}/2$, $-\ln(k_{F}a)=-1.5$, and $\kappa_{F}$ is the compressibility of a Fermi gas with the same SOC and total particle number but no pairing interaction.
@@
After showing that SOC and superfluidity are crucial for the singular behavior of $\kappa_{T}$ for the spherical case, we also verify the importance of the compact spherical geometry and the associated orbital angular momentum when the fermions are placed on it. In Appendix B, we present the results of $\Delta$, $\mu$, and $\kappa_{T}$ of Fermi superfluids on a 2D plane with a selected type of SOC, which may be considered as the flat-geometry limit of the spherical case. There is no sharp peak in $\kappa_{T}$ for the planar case, in stark contrast to the spherical results. The comparison between the 2D planar and spherical cases suggests that the singular behavior on a sphere arises from a combination of superfluidity, SOC, and geometry.
@@ V Implications
Since atomic BEC in a thin spherical shell has been realized using bubble traps or multi-species phase separation [1, 2, 3, 4, 5], similar techniques may apply to fermionic atoms to investigate Fermi gases on a spherical surface. As explained in Ref. [25], it may take more internal states and confining potentials to trap two-component Fermi gases in a thin spherical shell. After introducing tunable attractive interactions between the fermionic atoms via Feshbach resonance [56, 57], the ground state is expected to be a Fermi superfluid due to formation of Cooper pairs. Moreover, since a sphere has periodic boundary condition along any direction due to the spherical symmetry, the uniform solution with a constant order parameter $\Delta$ considered here is a natural choice. However, the spherical symmetry may be explicitly broken by imperfections or fluctuations in experiments. To investigate inhomogeneous effects without the spherical symmetry, one may resort to numerical methods like the Bogoliubov-de Gennes equation [26] or Hartree-Fock approximation [58] to explore possible structures of Fermi gases in the presence of SOC on a spherical surface.
@@
We mention that the presence of SOC can lead to very different behavior between Fermi gases with and without pairing interaction in a thin spherical shell by contrasting the aforementioned results and those in Appendix A. Instead of destroying the superfluid, SOC modifies the particle-hole mixed excitation spectrum and causes singular behavior in the free energy, which can be inferred by the kinks of the compressibility. Even without SOC, the compressibility of Fermi superfluid in 3D drastically differs from that on a spherical surface due to different dimensionality and geometry. Therefore, atomic Fermi superfluid in a thin spherical shell provides a playground for integrating superfluid, SOC, and geometric effects.
@@ Acknowledgements.
Y. H. was supported by the National Key R&\&&D Program of China (No. 2024YFF0508503). C. C. C. was partly supported by the NSF (No. PHY-2310656) and DOE (No. DE-SC0025809).
@@ Appendix A Fermi gases with only SOC on sphere
Therefore, if $\gamma^{2}>N$, there is no positive-$\mu$ solution.
For $\mu<0$, on the other hand, we always have $E_{1}^{(0)}>0$. The positive roots of the equation $E_{2}^{(0)}=0$ are $L^{(2)}_{1,2}=\dfrac{\gamma\pm\sqrt{\gamma^{2}+4\mu}}{2}$. This situation is shown in the bottom of Fig. 7. Thus, we find that
@@
Consequently, we find a solution $\mu=\dfrac{N^{2}-\gamma^{4}}{4\gamma^{2}}<0$ if $\gamma^{2}>N$. In summary, when $\gamma^{2}>N$, we have $\mu<0$ for the Fermi gas with only SOC. For a Fermi superfluid in the BCS limit, there will be a small $\Delta$, but we still expect that $\mu<0$ in this case.
@@
The corresponding compressibility is shown by the dotted line in Fig. 4. There is no singular behavior in the compressibility without pairing interaction as the strength of SOC varies. The difference between the Fermi gases with and without pairing interaction shows the importance of particle-hole mixing in the energy dispersions, which gives rise to interesting phenomena in superfluid only.
Appendix B 2D planar Fermi superfluid with SOC
Fermi superfluids in three-dimensional Euclidean space with various types of spin-orbital couplings have been discussed in Ref. [61], and no singular behavior of the compressibility was reported. Here we consider Rashba-type SOC on a 2D plane with the Hamiltonian
$H_{SOC}=\gamma(k_{x}\sigma_{y}-k_{y}\sigma_{x}).$ (59)
Here $k_{x}$ and $k_{y}$ are 2D the wave vectors and $\sigma_{a}$ with $a=x,y,z$ are the three Pauli matrices. This SOC term may be thought of as the planar limit of the spherical SOC term with the form $(\mathbf{r}\times\mathbf{k})\cdot\mathbf{S}$ with $S_{a}=\sigma_{a}/2$. The Hamiltonian of a 2D planar Fermi superfluid with Rashba-type SOC in momentum space can be written as
$H=\frac{1}{2}\sum_{\mathbf{k}}\Psi_{\mathbf{k}}^{\dagger}\left(\begin{array}[]{cccc}\xi_{\mathbf{k}}&\gamma(-k_{y}-ik_{x})&0&-\Delta\\ \gamma(-k_{y}+ik_{x})&\xi_{\mathbf{k}}-h&\Delta&0\\ 0&\Delta&-\xi_{\mathbf{k}}&\gamma(k_{y}-ik_{x})\\ -\Delta&0&\gamma(k_{y}+ik_{x})&-\xi_{\mathbf{k}}\end{array}\right)\Psi_{\mathbf{k}}+\sum_{\mathbf{k}}\xi_{\mathbf{k}}.$ (64)
Here $\Psi_{\mathbf{k}}=(c_{\mathbf{k},\uparrow},c_{\mathbf{k},\downarrow},c^{\dagger}_{-\mathbf{k},\uparrow},c^{\dagger}_{-\mathbf{k},\downarrow})^{T}$ and $\xi_{\mathbf{k}}=k^{2}/(2M)-\mu$ with $\mathbf{k}=(k_{x},k_{y})$.
Figure 8: Normalized chemical potential $\mu$ (left), order parameter $\Delta$ (right), and compressibility $\kappa_{T}$ as functions of $-\ln(k_{F}a)$ of 2D planar Fermi superfluids with Rashba-type SOC at $T=0$. The solid, dash, and dot-dash lines correspond to $\gamma=0$, $3$, and $5$, respectively.
For a given $\mathbf{k}$, the above 4 by 4 matrix can be diagonalized to give rise to four eigenvalues ±$\pm E_{1}(\mathbf{k})$ and ±$\pm E_{2}(\mathbf{k})$. Thus, there are two independent quasi-particle dispersions:
$E_{1}=\sqrt{\epsilon_{1}(\mathbf{k})^{2}+\Delta^{2}},\leavevmode\nobreak\ E_{2}=\sqrt{\epsilon_{2}(\mathbf{k})^{2}+\Delta^{2}},$ (65)
Here $\epsilon_{1}(\mathbf{k})=\xi_{\mathbf{k}}+\gamma\sqrt{k_{x}^{2}+k_{y}^{2}}$ and $\epsilon_{2}(\mathbf{k})=\xi_{\mathbf{k}}-\gamma\sqrt{k_{x}^{2}+k_{y}^{2}}$. The gap and number equations of the 2D Fermi superfluid with Rashba-type SOC at $T=0$ are given by
$\frac{1}{g}=\frac{1}{4}\sum_{\mathbf{k}}\Big{(}\frac{1}{E_{1}}+\frac{1}{E_{2}}\Big{)},$ (66)
$N=\sum_{\mathbf{k}}\Big{[}1-\frac{1}{2}\frac{\epsilon_{1}(\mathbf{k})}{E_{1}}-\frac{1}{2}\frac{\epsilon_{2}(\mathbf{k})}{E_{2}}\Big{]}.$ (67)
For a special case without SOC by setting $\gamma=0$, the two quasi-particle dispersions become $E_{1}=E_{2}=E_{\mathbf{k}}=\sqrt{\xi_{l}^{2}+\Delta^{2}}$, which recover the conventional results without SOC. Moreover, the gap equation can be regularized by the method discussed above Eq. (24). The compressibility is defined in Eq. (25). Following similar steps as before, the derivative $\partial N/\partial\mu$ can be computed as
$\frac{\partial N}{\partial\mu}=\Delta^{2}\mathcal{B}+\frac{\mathcal{A}^{2}}{\mathcal{B}}.$ (68)
Here $\mathcal{A}=\sum_{\mathbf{k}}\Big{[}\frac{\epsilon_{1}(\mathbf{k})}{2E_{1}^{3}}+\frac{\epsilon_{2}(\mathbf{k})}{2E_{2}^{3}}\Big{]}$ and $\mathcal{B}=\sum_{\mathbf{k}}\Big{(}\frac{1}{2E_{l}^{3}}+\frac{1}{2E_{2}^{3}}\Big{)}$.
Figure 8 shows $\mu$, $\Delta$, and $\kappa_{T}$ as functions of $-\ln(k_{F}a)$ at $T=0$ for $\gamma=0$, $3$, and $5$, respectively. Note that $\kappa_{T}$ is normalized by $\kappa_{F}=1/(2\pi)$, which is the compressibility of a 2D noninteracting Fermi gas. The curves of $\mu$ and $\Delta$ have similar trends as the spherical case with isotropic SOC shown in Fig. 1, but there is no kink in the 2D planar case. Furthermore, there is no sharp peak in $\kappa_{T}$ for the planar case. The contrast suggests that the sharp peaks in $\kappa_{T}$ of the spherical case has their origin in the compact geometry and the associated orbital angular momentum.